OpenStudy (anonymous):

Can anyone help me with this ? Which of the following options is an equivalent function to f(x) = 2(5)2x? f(x) = 50x f(x) = 100x f(x) = 2(25)x f(x) = 4(25)x

OpenStudy (anonymous):

What are the steps to solve this ?

rishavraj (rishavraj):

is it \[2(5)^2x\]????????

OpenStudy (anonymous):

yes @rishavraj

rishavraj (rishavraj):

whts the value of \[5^2 \] ???

OpenStudy (anonymous):

25

rishavraj (rishavraj):

and whts 25*2 ??

OpenStudy (anonymous):

50

rishavraj (rishavraj):

u r done then 50x

OpenStudy (anonymous):

thanks so much !

OpenStudy (anonymous):

Could you help me with another one ?

rishavraj (rishavraj):

yupppp

OpenStudy (anonymous):

For f(x) = 0.01(2)x, find the average rate of change from x = 3 to x = 8. 0.08 0.496 2.48 5

rishavraj (rishavraj):

is x as exponent ..????i mean is it \[0.01(2)x~~~or~~~0.1(2)^x\]

OpenStudy (anonymous):

It's 0.1(2) ^x

rishavraj (rishavraj):

avg = \[\frac{ 0.1(2)^8 - 0.1(2)^3 }{ 8 - 3 }\]

rishavraj (rishavraj):

avg rate of change

OpenStudy (anonymous):

4.96?

rishavraj (rishavraj):

oops it supposed to be 0.01

OpenStudy (anonymous):

The answer or something in the problem ?

rishavraj (rishavraj):

its supposed to be \[\frac{ 0.01 (2^8) - 0.01 (2^3)}{ 5 }\]

OpenStudy (anonymous):

i got 0.496

rishavraj (rishavraj):

is it in options??

OpenStudy (anonymous):

yes , is it the correct answer?

rishavraj (rishavraj):

yup :))) go fo r it

OpenStudy (anonymous):

okay thanks :)

rishavraj (rishavraj):

u r most welcome :))